Suppose you know the base of a rectangle has a length of 5 inches and a diagonal has a length of 13 inches. Find the area of the rectangle.
Found 2 solutions by ilana, linh117:
Answer by ilana(307)

If you sketch this rectangle, you should see its diagonal splits it into 2 right triangles. So the hypotenuse of the right triangle is 13 inches and one side is 5 inches. So the triangles are 5-12-13 triangles, and the third side (the other side of the rectangle) is 12 inches long. (You can also find this by using the Pythagorean Theorem.) So the area of the rectangle is the product of its length and width, or 5 x 12, or 60 square inches.

Answer by linh117(31)

solve like 2 triangles
A= 1/2 bh
Pythagorean theorem for third side
A^2 + B^2 = C^2
5^2 + B^2 = 13^2
25 + B^2 = 169
B^2 = 144
b= 12
A= 1/2(5)(12)
A=30
area of rectangle equal 30*2=60 inches^2